OpenStudy (anonymous):

Which is the equation of a circle with center (-2 3) and radius r=5? (x + 2)² + (y – 3)² = 10 (x + 2)² + (y – 3)² = 25 (x – 2)² + (y + 3)² = 10 (x – 2)² + (y + 3)² = 25

OpenStudy (anonymous):

4. Which is the equation of a circle with diameterwith A(5, 4) and B(–1, –4)? (x – 5)² + (y – 4)² = 10 (x + 5)² + (y + 4)² = 100 (x – 2)² + y² = 25 (x + 2)² + y² = 5

OpenStudy (anonymous):

2. A circle with center (–1, 2) passes through point (2, –2). Which is true? The radius is. sqrt 5 The diameter is 10. The equation is (x + 1)² + (y – 2)² = 10. The circumference is 25.

OpenStudy (anonymous):

for your 1st question the answer is B for the 3rd question the answer is the diameter is 10 \[r^2=(x-h)^2+(y-k)^2\rightarrow (2+1)^2+(-2-2)^2\] which you have \[r^2=3^2+4^2\rightarrow r^2=25\rightarrow r=5\] diameter is 2*r so \[2(5)=10\] i'm currently working on the 2nd one

OpenStudy (anonymous):

the answer for the 2nd one is B

OpenStudy (anonymous):

number 4 was C not B